OpenStudy (anonymous):

Currently, Sam's coach is three times older than Sam. When Sam started soccer 5 years ago, his coach was five times his age. How old is Sam now?

OpenStudy (mertsj):

S = Sam's age 3S= Coach's age S-5 = Sam's age 5 years ago 3S-5 = Coach's age 5 years ago. 5(S-5)=3S-5